OpenStudy (anonymous):

Radio signals travel at a rate of 3.00 × 10^8 meters per second. How many seconds would it take for a radio signal to travel from a satellite to the surface of the Earth if the satellite is orbiting at a height of 3.54 * 10^7 @johnweldon1993 (also my computers about to die so could we do this as fast as possible? THANK YOU SOOO MUCH)

OpenStudy (anonymous):

I have to show my work @johnweldon1993

OpenStudy (johnweldon1993):

\[\large \frac{3.54 \times 10^7}{3.00 \times 10^8} = ?\]

OpenStudy (anonymous):

:/ Still confused.. Do I convert to regular notion?

OpenStudy (anonymous):

I'm having bad internet so if I already asked that i'm sorry